A service provider measures availability over a month by dividing it into 10-minute intervals and performing a synthetic transaction testevery interval.  If the synthetic transaction test fails, the system is considered "down" for the entire 10 minute interval.  Scheduled downtime is excluded from agreed service time.

In the month of January, the service provider took one 7 hour window of scheduled downtime, and the synthetic transaction test failed 21 times outside of the scheduled downtime window.  What was the availability of the system for that month?

Express your answer as a percentage (so if the answer is 12.34% then enter the value 12.34).
Answer #1

Total Available Hours

= 31 Days *24 hours *60 minutes

=44640 minutes

Scheduled downtine

=7 hours*60 minutes

=420 minutes

Total service time

= 44640-420= 44220 minutes

Unscheduled down time

=21 times * 10 minutes = 210 minutes

Total Availability

=44220-210= 44010 minutes

Total Availability in percentage

=44010/44220

=99.525%

Final answer is 99.525%